Joy invests a total of $33,000 in two accounts paying 9% and 8% annual interest, respectively. How much was invested in each account if, after one year, the total interest was $2,780.00.
How much was invested at 9% and how much was invested at 8%?

Answer by mananth(16946)

Part I 9.00% per annum ------------- Amount invested =x
Part II 8.00% per annum ------------ Amount invested = y
33000
Interest----- 2780.00

Part I 9.00% per annum ---x
Part II 8.00% per annum ---y
Total investment
x + 1 y= 33000 -------------1
Interest on both investments
9.00% x + 8.00% y= 2780
Multiply by 100
9 x + 8 y= 278000.00 --------2
Multiply (1) by -9
we get
-9 x -9 y= -297000.00
Add this to (2)
0 x -1 y= -19000
divide by -1
y = 19000
Part I 9.00% $ 14000
Part II 8.00% $ 19000

CHECK
14000 --------- 9.00% ------- 1260.00
19000 ------------- 8.00% ------- 1520.00
Total -------------------- 2780.00

m.ananth@hotmail.ca